OpenStudy (anonymous):

rewrite the expression as an algebraic function of x and y sin(cos^(-1)(x)tan^(-1)(y))

OpenStudy (anonymous):

\[\sin(\cos^{-1} (x)\tan^{-1} (y))\]

OpenStudy (anonymous):

maybe \[\sin(\cos^{-1} (x)+\tan^{-1} (y))\]?? otherwise i don't think you are going to get an expression in x